# Unbiased Sliced Wasserstein Kernels for High-Quality Audio Captioning

### v1

@@ 5.3 Qualitative Experiments
Human evaluation. We conduct a human evaluation to better assess the quality of generated captions. We randomly choose 50 audio from AudioCaps and Clotho test data. Captions are generated for each audio by using different methods: maximum likelihood estimation (MLE), contrastive framework, and the ACUS framework. The MLE method utilizes a deterministic decoding method, beam search with a beam size of 5, while contrastive learning and the proposed method utilize a stochastic decoding method, top-p sampling with $p=0.7$ to generate 30 candidate captions. The most suitable caption is chosen based on Equation (5) for contrastive learning and Equation (14) for the proposed method. We recruit five annotators, who are asked to independently assess the quality of a given caption following a 5-point Likert scale for three aspects
  • Descriptiveness: Whether the caption is descriptive enough, describe all audio events in the given audio and their temporal relationships.
  • Correctness: Whether the caption is correct, all audio events occur in the given audio.
  • Fluency: Whether the caption is fluent and easy to understand as human written.
Table 4 shows the human valuation results on three aspects for Audiocaps and Clotho datasets. The inter-annotator agreement is shown in the last row measured by the Fleiss Kappa score (Fleiss, 1971). On both datasets, our method is capable of generating more descriptive and correct captions compared to baseline models trained with MLE and contrastive learning objectives. Also, all generated captions are more fluent than human-written captions. The rationale behind it is that humans focus more on audio content rather than fluency. On the other hand, audio captioning models leverage pretrained language models as the decoder, therefore, they can generate coherence captions but less focus on describing audio content. The qualitative examples can be found in Appendix A.4.
5.4 Ablation Studies
@@
Table 5 shows the ablation study on choosing similarity metrics for measuring audio and caption similarity. The DTW and soft- DTW are ineffective in measuring the similarity across acoustic and linguistic modality. Therefore, there is a decrease in performance compared with the baseline method with beam search decoding. The hypothesis is that the constraint for monotonic alignment between acoustic and linguistic embedding is too strict for measuring the distance between two modalities. Our score and the Wasserstein distance relax the monotonic alignment constraint when computing cross-modality similarity. Both our score and the Wasserstein distance are equipped with the positional embedding to consider temporal information when measuring similarity across modalities. Relaxing the monotonic alignment and incorporating positional embedding(PE) shows a significant performance gain regarding METEOR and SPICE metrics with the Wasserstein distance, $0.254$ to $0.262$ and $0.186$ to $0.194$, respectively. Although the Wasserstein distance with positional embedding is effective in measuring acoustic and linguistic similarity, it possesses a weakness: the dimensionality curse. Thus, there is still a gap in calculating similarity across acoustic and linguistic modalities. As mentioned in (Nguyen & Ho, 2022; Nietert et al., 2022; Nadjahi et al., 2020), the sliced Wasserstein does not suffer from the dimensionality curse. The performance of the USW-RBF score acquires a performance gain with all evaluation metrics, which reflects that the sliced Wasserstein with positional embedding is the most effective score for computing audio and caption similarity. The ablation study on the number of Monte Carlo samples $L$ for estimating the USW-RBF is shown in Table 9 in Appendix A.3.
Method Inference time on a A6000 GPU(RTF)
MLE 0.33
MLE + CL 0.65
MLE + ACUS 0.81
@@
We conducted an ablation study on the effectiveness of positional embedding techniques for our method. As shown in Table 6, the rotary positional embedding technique outperforms the absolute positional embedding technique regarding all evaluation metrics. The rotary positional embedding (PE) technique outperforms both without PE and the absolute PE technique regarding all objective metrics. These empirical results indicate that the rotary PE technique is the most suitable method for the ACUS framework to account for temporal information when measuring cross-modal similarity. We also demonstrated the real-time-factor of our proposed framework in Table. 7. Although the inference time of the ACUS framework is the longest, it is still able to generate audio captions in real-time. Therefore, it can be deployed for real-world applications.
@@ 6 Conclusion
We introduce the ACUS framework for alleviating text degeneration for the audio captioning task. Furthermore, we develop the USW-RBF kernel equipped with the rotary positional embedding. The USW-RBF is an unbiased kernel, thus, it is compatible with stochastic gradient optimization algorithms, and its approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$. Our experiments demonstrate that our framework is able to mitigate the text degeneration issue for audio captioning models and outperforms baseline methods in terms of quantitative and qualitative evaluations. We further find that the nucleus sampling technique is the best decoding method to generate descriptive and correct captions from pretrained audio captioning models.
@@ A.1 Proofs
A.1.2 Proof of Proposition 3.3
We first recall the definition of SW-RBF (Equation (8)) and the definition of USW-RBF (Definition 3.1.
@@ A.2 Implementation details
where $Z_{X}\in\mathbb{R}^{n\times d}$ and $Z_{y}\in\mathbb{R}^{m\times d}$ are two $d-$dimensional sequences of audio and text hidden representation. The cost matric between them is denoted as $C(Z_{X},Z_{Y})$, in which its element is computed as $c_{i,j}=\frac{1}{2}||z_{x}^{i}-z_{y}^{j}||^{2}_{2}$. We denote $\mathcal{A}(m,n)\subset{0,1}^{m\times n}$ as a set of all such monotonic alignment matrices. The soft- DTW is a variant of  DTW which is compute as follow
@@
A.3 Ablation studies

### v2

@@ 5.4 Ablation study
Table 6 shows the ablation study on choosing similarity metrics for measuring audio and caption similarity. The DTW and soft- DTW are ineffective in measuring the similarity across acoustic and linguistic modality. Therefore, there is a decrease in performance compared with the baseline method with beam search decoding. The hypothesis is that the constraint for monotonic alignment between acoustic and linguistic embedding is too strict for measuring the distance between two modalities. Our score and the Wasserstein distance relax the monotonic alignment constraint when computing cross-modality similarity. Both our score and the Wasserstein distance are equipped with the positional embedding to consider temporal information when measuring similarity across modalities. Relaxing the monotonic alignment and incorporating positional embedding(PE) shows a significant performance gain regarding METEOR and SPICE metrics with the Wasserstein distance, $0.254$ to $0.262$ and $0.186$ to $0.194$, respectively. Although the Wasserstein distance with positional embedding is effective in measuring acoustic and linguistic similarity, it possesses a weakness: the dimensionality curse. Thus, there is still a gap in calculating similarity across acoustic and linguistic modalities. As mentioned in (Nguyen and Ho, 2022; Nietert et al., 2022; Nadjahi et al., 2020), the sliced Wasserstein does not suffer from the dimensionality curse. The performance of the USW-RBF score acquires a performance gain with all evaluation metrics, which reflects that the sliced Wasserstein with positional embedding is the most effective score for computing audio and caption similarity.
@@
We conducted an ablation study on the effectiveness of positional embedding techniques for our method. As shown in Table 7, the rotary positional embedding technique outperforms the absolute positional embedding technique regarding all evaluation metrics. The rotary positional embedding (PE) technique outperforms both without PE and the absolute PE technique regarding all objective metrics. These empirical results indicate that the rotary PE technique is the most suitable method for the ACUS framework to account for temporal information when measuring cross-modal similarity. We also conducted an ablation study on the inference time in appendix. A.4.
@@ 6 Conclusion
We introduce the ACUS framework for alleviating text degeneration for the audio captioning task. Furthermore, we develop the USW-RBF kernel equipped with the rotary positional embedding. The USW-RBF is an unbiased kernel, thus, it is compatible with stochastic gradient optimization algorithms, and its approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$. Our experiments demonstrate that our framework is able to mitigate the text degeneration issue for audio captioning models and outperforms baseline methods in terms of quantitative and qualitative evaluations. We further find that the nucleus sampling technique is the best decoding method to generate descriptive and correct captions from pretrained audio captioning models. The experiments on audio reasoning tasks also demonstrate the generalizability of our kernel on a broad range of cross-modal audio-text tasks.
Acknowledgments and Disclosure of Funding
Dinh Phung is supported by the Australian Research Council (ARC) Discovery Project DP250100262 and DP230101176.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims are supported by the analysis and extensive experiments in Section 3 and Section 5, respectively.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of our proposed method in Appendix. A4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide proofs for our theory in the Appendix. A1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The implementation details are described in the Appendix A2 for reproducibility purposes.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The source code for experiments is uploaded in the supplementary. The GitHub link and pretrained models will be released when the manuscript gets accepted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: all experimental settings are described in the experiment section and the Appendix A2.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: all experiments related to subjective evaluation and stochastic sampling are reported with appropriate error bars to show statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: all experimental settings are provided in the experiment section and the Appendix A2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [N/A]
  44. Justification: N/A
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: N/A
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: N/A
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: N/A
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: N/A
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: details regarding subjective evaluation are described in Section 5.2.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: N/A
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: N/A
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ A.1.1 Proof of Proposition 1
A.1.2 Proof of Proposition 2
We first recall the definition of SW-RBF (Equation (8)) and the definition of USW-RBF (Definition 1.
@@ A.2 Implementation details
Baselines. We compare against all state-of-the-art audio captioning models on the Audiocaps and Clotho datasets. The AutoCap model Haji-Ali et al. (2024) leverages a compact representation from the CLAP encoders and audio metadata to enhance audio caption quality. LHDFF (Sun et al., 2023) utilizes residual the PANNs encoder to fuse low and high dimensional features in Mel-spectrogram. CNN14-GPT2 (Kim et al., 2023) and Pengi (Deshmukh et al., 2023) apply prefix-tuning method for the pretrained GPT2 (Radford et al., 2019). The BART-tags (Gontier et al., 2021) model generates audio captions relying on predefined audio tags from the AudioSet dataset. AL-MixGen (Kim et al., 2022) leverages the ACT backbone trained using audio-language mixup augmentation and test-time augmentation at the inference phase. Wavcaps Mei et al. (2024) is the HTSAT-BART model Chen et al. (2022b) fine-tuned on numerous weakly-labeled data which is generated by using large language models. We choose a subset of models evaluated on the Clotho dataset without complex training methods, such as ensemble training, to ensure a fair comparison. The CLIP-AAC (Chen et al., 2022a), MAAC (Ye et al., 2021), P-LocalAFT(Xiao et al., 2022), and Graph-AC (Xiao et al., 2023) are the baselines evaluated on Clotho dataset.
ACT backbone. We follow the original settings in (Mei et al., 2021) to train the audio captioning transformer (ACT) backbone on the AudioCaps dataset. We use a batch size of 32 and warm up for five epochs before reaching the peak learning rate at $lr=1e^{-4}$. We use the training objective function in Equation (13) with training coefficient $\alpha=0.1$ and the bandwidth for the temporal-similarity metric $\gamma=2.5$. We also perform a grid search for stochastic decoding methods at the inference state to choose the best hyperparameters for each stochastic decoding method, $p=\{0.5,0.6,0.7,0.8,0.9\}$ for top-p sampling, $k=\{3,4,5\}$ for top-k sampling, and $\text{temp}=\{1.1,1.2,1.3,1.4,1.5\}$ for temperature sampling. The best results with optimal decoding hyperparameters are reported in Table 11.
@@
[…]
@@
A.3 Additional ablation studies
Table 8: Ablation study on the effectiveness of the proposed USW-RBF kernel on the AudioCaps dataset with the Enclap backbone. Both baseline Enclap and the baseline Enclap with the USW-RBF kernel in training utilize a deterministic decoding technique (beam search with $\text{beam size}=5$). The decoding method is top-p sampling with $p=0.7$ for the ACUS framework.
PE method METEOR ROUGE_L CIDEr SPICE SPIDEr
Enclap 0.254 0.5 0.77 0.186 0.48
Enclap + USW-RBF in training 0.256 0.496 0.79 0.188 0.492
Enclap + USW-RBF in both (ACUS) $\textbf{0.262}\pm\textbf{0.001}$ $\textbf{0.509}\pm\textbf{0.001}$ $\textbf{0.807}\pm\textbf{0.003}$ $\textbf{0.193}\pm\textbf{0.001}$ $\textbf{0.5}\pm\textbf{0.002}$